OpenStudy (anonymous):

Which is an expression for the derivative P(x)=1+f(x)/f(x) , where f(x) is a differentiable function?

OpenStudy (anonymous):

\[P(x)=\frac{f(x)+1}{f(x)}\] ?

OpenStudy (anonymous):

yes. :)

OpenStudy (anonymous):

\[P(x)=1+\frac{1}{f(x)}\] \[P'(x)=\frac{-f'(x)}{f^2(x)}\]

OpenStudy (anonymous):

Re-write it as,\[P(x)=1+\frac{1}{f(x)}\]We need to use the chain rule here:\[dP/dx=\frac{dP}{df(x)}\frac{df(x)}{dx}\]This gives:\[dP/dx=\frac{-1}{f(x)^2}f \prime(x)\]

hero (hero):

bravo

OpenStudy (anonymous):

I wrote the solution a bit confusing, I mean it as:\[dP/dx=\frac{-1}{[f(x)]^2}f \prime(x)\]

OpenStudy (anonymous):

you're the best!